# New Physics at Tera-$Z$: Precision Renormalised

### v1

@@ Appendix A EWPO operators generated by each state
$\mathcal{S}$ $\mathcal{O}_{H\Box}$, $\mathcal{O}_{HB}$, $\mathcal{O}_{HW}$ ($\mathcal{O}^{(1)}_{Hq}$, $\mathcal{O}^{(3)}_{Hq}$, $\mathcal{O}^{(1)}_{Hl}$, $\mathcal{O}^{(3)}_{Hl}$, $\mathcal{O}_{He}$)${}^{g^{2},y^{2}}$,
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Xi}$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}_{HD}}$, $\mathcal{O}_{H\Box}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}_{HWB}}$
@@
Table 6: Dimension-six operators contributing to EWPOs generated by the heavy scalar fields introduced in Ref. deBlas:2017xtg . Operators and models entering EWPOs at tree-level are shown in red. Operators which upon RG evolution generate EWPO operators are shown in black. In blue are operators generated at one-loop at the matching scale Durieux:2022hbu . The last column details the EWPO operators generated at one-loop through RG evolution. For models already constrained at tree-level the one-loop contributions are not shown. The superscript denotes the order in couplings $g$ for EW gauge, $y$ for the associated Yukawa coupling, at which the Wilson coefficient arises.
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U}$ $\mathcal{O}_{uB}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}^{(1)}_{Hq}}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}^{(3)}_{Hq}}$
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Q_{1}}$ $\mathcal{O}_{uB}$, $\mathcal{O}_{uW}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}_{Hd}}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}_{Hu}}$
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}T_{2}}$ $\mathcal{O}_{uW}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}^{(1)}_{Hq}}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}^{(3)}_{Hq}}$

### v2

@@ Appendix A EWPO operators generated by each state
$\mathcal{S}$ $\mathcal{O}_{H\Box}$, [$\mathcal{O}_{HB}$], [$\mathcal{O}_{HW}$] ($\mathcal{O}^{(1)}_{Hq}$, $\mathcal{O}^{(3)}_{Hq}$, $\mathcal{O}^{(1)}_{Hl}$, $\mathcal{O}^{(3)}_{Hl}$, $\mathcal{O}_{He}$)${}^{g^{2},y^{2}}$,
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Xi}$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}_{HD}}$, $\mathcal{O}_{H\Box}$, $[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}_{HWB}}$]
@@
Table 6: Dimension-six operators contributing to EWPOs generated by the heavy scalar fields introduced in Ref. deBlas:2017xtg . Operators enclosed within square brackets are only generated via non-renormalisable interactions of the given scalar. Operators and models entering EWPOs at tree-level are shown in red. Operators which upon RG evolution generate EWPO operators are shown in black. In blue are operators generated at one-loop at the matching scale Durieux:2022hbu . The last column details the EWPO operators generated at one-loop through RG evolution. For models already constrained at tree-level the one-loop contributions are not shown. The superscript denotes the order in couplings $g$ for EW gauge, $y$ for the associated Yukawa coupling, at which the Wilson coefficient arises.
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}U}$ [$\mathcal{O}_{uB}$], ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}^{(1)}_{Hq}}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}^{(3)}_{Hq}}$
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Q_{1}}$ [$\mathcal{O}_{uB}$], [$\mathcal{O}_{uW}$], ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}_{Hd}}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}_{Hu}}$
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}T_{2}}$ [$\mathcal{O}_{uW}$], ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}^{(1)}_{Hq}}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathcal{O}^{(3)}_{Hq}}$